Jacqueline A. & Sean B. Company purchases goods from its suppliers on terms of 4/10, net 40. (Show Calculations)

a. What is the effective annual cost to Jacqueline & Sean Company if it chooses not to take the discount and makes its payment on day 40?

b. What is the effective annual cost to Jacqueline & Sean Company if it chooses not to take the discount and makes it payment on day 60?
Answer #1

The effective annual cost in the two situations can be calculated as follows:

Effective Annual Cost = (1+ (Discount% / (100 - Discount%)))^(365/(Days To Pay - Discount Period)) - 1

Part (a)

Discount = 4, Days To Pay = 40 and Discount Period = 10
Using these values in the above formula, we get,
Effective Annual Cost = (1+(4/(100-4)))^(365/(40 - 10)) - 1 = 64.32%

Part (b)

Discount = 4, Days To Pay = 60 and Discount Period = 10
Using these values in the above formula, we get,
Effective Annual Cost = (1+(4/(100-4)))^(365/(60 - 10)) - 1 = 34.72%


answered by: studyres